OpenStudy (anonymous):

My sister needs help: What is... -4+7(1-3m)= ? -5n+3(6+7n)= ?

OpenStudy (calculusxy):

You first need to distribute the term outside of the parenthesis to the terms inside of the parenthesis. \(x(a + b) = x(a) + x(b)\) Similarly, for question 1, what you do is: \(-4 + 7(1 - 3m) = -4 + 7(1) + 7(-3m)\) \(4 + 7 - 21m\) combine the like terms to get \(11 - 21m\). Use the same principle for the second question.

OpenStudy (anonymous):

Thanks @calculusxy

OpenStudy (calculusxy):

You're welcome :)